# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ Appendix A Proofs for Section 2
Hence, the loss for the tabular MDP problem satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]$. Using Lemma 5 with $n=1$, we can conclude the the loss for the tabular MDP problem satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$. From Mei et al. (2020, Lemma 8), we know that
$\left\|\nabla f(\theta)\right\|$ $\geq\frac{\min_{s}\pi_{\theta}(\pi^{*}(a)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}\rho(s)}}\,f(\theta)$
Hence, the loss for the tabular MDP problem satisfies Assumption 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(\pi^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$. ∎
@@
Using the proof of Proposition 1 to calculate the gradient and hessian, we get that $\nabla f_{1}(0)=0$ and $\nabla f_{2}(0)=0$ which implies $\nabla f(0)=0$. Similarly, for Hessian, we get $\nabla^{2}f_{1}(0)=1$ and $\nabla^{2}f_{2}(0)=1$ which implies $\nabla^{2}f(0)=1$. Since $\nabla f(\theta)=0$ and $\nabla^{2}f(\theta)\neq 0$, the assumption cannot hold with $L_{0}\neq 0$ and any $L_{1}\neq 0$. ∎
Appendix B Proofs for Section 3
@@
Case 1: If $L_{1}=0$Assumption 2 is equivalent to the standard $L_{0}$-uniform smoothness condition. In this case, we can follow the standard analysis of GD-LS (Nocedal & Wright, 2006) and conclude that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L_{0}}\right\}\geq\min\left\{\eta_{\max},\frac{(1-c)}{3\,L_{0}}\right\}$.
@@
Using Assumption 2, $\nabla^{2}f(\theta)\preceq[L_{0}+L_{1}\,f(\theta)]\,I_{d}$. Using this result, we bound the Hessian of $g(\theta)$.
@@
$\implies\ln\left(L_{0}+L_{1}\,f(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))\right)$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$
$\implies g(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$
$=\ln\left((L_{0}+L_{1}\,f(\theta_{t}))\,\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)\right)=g(\theta_{t})+\ln\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)$
@@
If the Armijo condition is satisfied for an ${\eta_{t}}$ s.t. $h_{Q}({\eta_{t}})\leq h_{L}({\eta_{t}})$, it implies that ${\eta_{t}}\leq\frac{2(1-c)}{L_{0}+L_{1}f(\theta_{t})}$. However, we show that the resulting step-size cannot be too small. In particular, we will prove that the Armijo condition is satisfied for ${\eta_{t}}=\frac{2(1-c)}{6(L_{0}+L_{1}\omega)+6\,L_{1}\,(\nu+1)\,f(\theta_{t})}$. To show this, we use Assumption 2. In order to use this inequality, we have to ensure that $\left\|\theta_{t+1}-\theta_{t}\right\|={\eta_{t}}\,\left\|\nabla f(\theta_{t})\right\|\leq\frac{q}{L_{1}}$. Since based on Assumption 3, $\left\|\nabla f(\theta_{t})\right\|\leq\nu\,f(\theta_{t})+\omega$, it suffices to ensure that $q\geq{\eta_{t}}\,L_{1}(\nu\,f(\theta_{t})+\omega)$.
Using Lemma 5, we get that:
@@
$\implies{\eta_{t}}$ $\leq\frac{2\,(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})}$
@@
Moreover, for ${\eta_{t}}^{\prime}:=\frac{2(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)}$, we need to ensure that $q\geq{\eta_{t}}^{\prime}\,L_{1}(\nu\,f(\theta_{t})+\omega)$. Hence, we want to find a $q$ s.t.
@@
Based on above argument, we can conclude that the ${\eta_{t}}$, the step-size returned by the Armijo line-search is lower-bounded as ${\eta_{t}}\geq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}$. Moreover if $\eta_{\max}\leq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}$, then $\eta_{\max}$ satisfies both Armijo condition and $h_{Q}(\eta_{\max})\leq h_{L}(\eta_{\max})$, in which case, the line-search would terminate immediately and return $\eta_{\max}$. Therefore ${\eta_{t}}\geq\min\{\eta_{\max},\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}\}$. ∎
@@
We now follow a proof similar to that of Axiotis & Sviridenko (2023, Theorem 5.2) and derive a linear rate of convergence. From the theorem assumption, we know that $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\frac{[f(\theta)-f^{*}]^{2}}{R}$. Combining these relations,
@@
Case (1): If ${f^{*}}\geq\frac{\lambda_{0}}{\lambda_{1}}-\epsilon$. Since $\{f(\theta_{t})\}_{t=0}^{t=\tau}$ is monotonically decreasing due to the Armijo line-search and converging to $\frac{\lambda_{0}}{\lambda_{1}}$. Hence, there exists a $\tau^{\prime}$ s.t. $\tau^{\prime}\leq\tau$ such that $f(\theta_{\tau^{\prime}})-{f^{*}}\leq\epsilon$ and $f(\theta_{\tau^{\prime}-1})-{f^{*}}\geq\epsilon$, i.e. $\tau^{\prime}$ is the iteration index when the desired sub-optimality criterion is satisfied for the first time. This implies that for all $t<\tau^{\prime}$, $\delta_{t}:=f(\theta_{t})-{f^{*}}>\epsilon$. Hence, $\frac{f(\theta_{\tau^{\prime}})}{{f^{*}}}\leq\gamma:=1+\frac{\epsilon}{{f^{*}}}$. Since ${f^{*}}>0$ and $\epsilon>0$, $\gamma>1$. Hence for all $t<\tau^{\prime}$,
$\frac{f(\theta_{t})}{{f^{*}}}>\gamma\implies\frac{\delta_{t}}{f(\theta_{t})}=1-\frac{{f^{*}}}{f(\theta_{t})}>1-\frac{1}{\gamma}>0.$
@@
$\delta_{t+1}$ $\leq\left(1-\underbrace{\bar{\alpha}\,\left(1-\frac{1}{\gamma}\right)}_{:=\rho}\right)\delta_{t}$
Since $\bar{\alpha}\in(0,1)$ and $\left(1-\frac{1}{\gamma}\right)\in(0,1)$, $\rho:=\bar{\alpha}\,\left(1-\frac{1}{\gamma}\right)\in(0,1)$. Recursing from $t=0$ to $t=\tau^{\prime}-1$,
@@ Appendix C Proofs for Section 4
Using Theorem 1 with $R=\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2}$ and setting $L_{0}=0$ completes the proof. ∎
@@
For an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
@@
$=\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,{\eta_{t}}\,[f(\theta_{t})-f(u)]$ (14)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,\frac{1}{\lambda_{0}+\lambda_{1}f(\theta_{t})}[f(\theta_{t})-f(u)]$ (Using Lemma 1)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-(2-\frac{1}{c})\,\frac{1}{\lambda_{1}}\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}$ ( using $\lambda_{0}=0$ since $L_{0},\omega=0$ and by defining $C:=(2-\frac{1}{c})\,(1/\lambda_{1})$)
@@
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\gamma:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\gamma$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\gamma}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\gamma}\right)$
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{f(u)}{\epsilon}\right]$ (Using the definition of $\gamma$)
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Theorem 2. In particular, for a comparator $u=\beta u^{*}$, GD with Armijo line-search with $c$, $\eta_{\max}=\infty$ and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{2c-1}{c\lambda_{1}}$,
@@
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{2}\right]$
@@
$=\frac{6\,c\,}{(2c-1)(1-c)\gamma^{2}}\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$ (using Proposition 1 for the value of $\nu$ and $L_{1}$)

### v2

@@ Acknowledgments
Hence, the loss for the tabular MDP problem satisfies Assn. 5 with $L_{c}=0$ and $L_{g}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\sqrt{S}$. Using Lemma 4 with $n=1$, we can conclude the the loss for the tabular MDP problem satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\,\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{2}\,S$ and Assn. 3 with $\nu=8\,\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\cdot\sqrt{S}$ and $\omega=0$. From Mei et al. (2020, Lemma 8), we know that
$\left\|\nabla f(\theta)\right\|$ $\geq\frac{\min_{s}\pi_{\theta}(a^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}\rho(s)}}\,f(\theta)$
Hence, the loss for the tabular MDP problem satisfies Assn. 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(a^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$. ∎
@@
Using the proof of Prop. 1 to calculate the gradient and hessian, we get that $\nabla f_{1}(0)=0$ and $\nabla f_{2}(0)=0$ which implies $\nabla f(0)=0$. Similarly, for Hessian, we get $\nabla^{2}f_{1}(0)=1$ and $\nabla^{2}f_{2}(0)=1$ which implies $\nabla^{2}f(0)=1$. Since $\nabla f(\theta)=0$ and $\nabla^{2}f(\theta)\neq 0$, the assumption cannot hold with $L_{0}\neq 0$ and any $L_{1}\neq 0$. ∎
Appendix B Proofs for Sec. 3
@@
Case 1: If $L_{1}=0$Assn. 2 is equivalent to the standard $L_{0}$-uniform smoothness condition. In this case, we can follow the standard analysis of GD-LS (Nocedal & Wright, 2006) and conclude that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L_{0}}\right\}\geq\min\left\{\eta_{\max},\frac{(1-c)}{3\,L_{0}}\right\}$.
@@
Using Assn. 2, $\nabla^{2}f(\theta)\preceq[L_{0}+L_{1}\,f(\theta)]\,I_{d}$. Using this result, we bound the Hessian of $g(\theta)$.
@@
Note that $L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\ \left\|\nabla f(\theta_{t})\right\|_{2}^{2}\geq 0$. Since $f,L_{0},L_{1}\geq 0$, $1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\geq 0$. Taking log on both sides,
$\ln\left(L_{0}+L_{1}\,f(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))\right)$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$ (Since $\ln$ is monotonically increasing)
$\implies g(\theta_{t}-{\eta_{t}}\nabla f(\theta_{t}))$ $\leq\ln\left(L_{0}+L_{1}\,f(\theta_{t})-c{\eta_{t}}\,L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\right)$ (By definition of $g$)
$=\ln\left((L_{0}+L_{1}\,f(\theta_{t}))\,\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)\right)$ (14)
$=g(\theta_{t})+\ln\left(1-c{\eta_{t}}\,\frac{L_{1}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}{L_{0}+L_{1}\,f(\theta_{t})}\right)$ (By definition of $g$)
@@
If the Armijo condition is satisfied for an ${\eta_{t}}$ s.t. $h_{Q}({\eta_{t}})\leq h_{L}({\eta_{t}})$, it implies that ${\eta_{t}}\leq\frac{2(1-c)}{L_{0}+L_{1}f(\theta_{t})}$.
However, we show that the resulting step-size cannot be too small. In particular, we will prove that the Armijo condition is satisfied for
${\eta_{t}}=\frac{2(1-c)}{6(L_{0}+L_{1}\omega)+6\,L_{1}\,(\nu+1)\,f(\theta_{t})}.$
To show this, we use Assn. 2. In order to use this inequality, we have to ensure that $\left\|\theta_{t+1}-\theta_{t}\right\|={\eta_{t}}\,\left\|\nabla f(\theta_{t})\right\|\leq\frac{q}{L_{1}}$. Since based on Assn. 3, $\left\|\nabla f(\theta_{t})\right\|\leq\nu\,f(\theta_{t})+\omega$, it suffices to ensure that $q\geq{\eta_{t}}\,L_{1}(\nu\,f(\theta_{t})+\omega)$.
Using Assn. 2, we get that:
@@
$\hskip 107.63855pt{\eta_{t}}\leq\frac{2\,(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})},$
@@
Moreover, for
${\eta_{t}}^{\prime}:=\frac{2(1-c)}{\left(1+e^{q}-\frac{e^{q}-1}{q}\right)\,L_{0}+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,f(\theta_{t})+\left(\frac{e^{q}-1}{q}\right)\,L_{1}\,(\nu f(\theta_{t})+\omega)},$
we need to ensure that $q\geq{\eta_{t}}^{\prime}\,L_{1}(\nu\,f(\theta_{t})+\omega)$. Hence, we want to find a $q$ s.t.
@@
Based on above argument, we can conclude that the ${\eta_{t}}$, the step-size returned by the Armijo line-search is lower-bounded as
${\eta_{t}}\geq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}.$
Moreover if
$\eta_{\max}\leq\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})},$
then $\eta_{\max}$ satisfies both Armijo condition and $h_{Q}(\eta_{\max})\leq h_{L}(\eta_{\max})$, in which case, the line-search would terminate immediately and return $\eta_{\max}$. Therefore
${\eta_{t}}\geq\min\{\eta_{\max},\frac{2(1-c)}{6\,(L_{0}+L_{1}\omega)+6\,L_{1}(\nu+1)\,f(\theta_{t})}\}.$
@@
We now follow a proof similar to that of Axiotis & Sviridenko (2023, Theorem 5.2) and derive a linear rate of convergence. From the theorem assumption, we know that $\left\|\nabla f({\theta_{t}})\right\|_{2}^{2}\geq\frac{[f({\theta_{t}})-f^{*}]^{2}}{R}$. Combining these relations,
@@
Case (1): If ${f^{*}}\geq\frac{\lambda_{0}}{\lambda_{1}}-\epsilon$. Since $\{f(\theta_{t})\}_{t=0}^{t=\tau}$ is monotonically decreasing due to the Armijo line-search and converging to $\frac{\lambda_{0}}{\lambda_{1}}$. Hence, there exists a $\tau^{\prime}$ s.t. $\tau^{\prime}\leq\tau$ such that $f(\theta_{\tau^{\prime}})-{f^{*}}\leq\epsilon$ and $f(\theta_{\tau^{\prime}-1})-{f^{*}}\geq\epsilon$, i.e. $\tau^{\prime}$ is the iteration index when the desired sub-optimality criterion is satisfied for the first time. This implies that for all $t<\tau^{\prime}$, $\delta_{t}:=f(\theta_{t})-{f^{*}}>\epsilon$. Hence, $\frac{f(\theta_{\tau^{\prime}})}{{f^{*}}}\leq\alpha:=1+\frac{\epsilon}{{f^{*}}}$. Since ${f^{*}}>0$ and $\epsilon>0$, $\alpha>1$. Hence for all $t<\tau^{\prime}$,
$\frac{f(\theta_{t})}{{f^{*}}}>\alpha\implies\frac{\delta_{t}}{f(\theta_{t})}=1-\frac{{f^{*}}}{f(\theta_{t})}>1-\frac{1}{\alpha}>0.$
@@
$\delta_{t+1}$ $\leq\left(1-\underbrace{\bar{\alpha}\,\left(1-\frac{1}{\alpha}\right)}_{:=\rho}\right)\delta_{t}$
Since $\bar{\alpha}\in(0,1)$ and $\left(1-\frac{1}{\alpha}\right)\in(0,1)$, $\rho:=\bar{\alpha}\,\left(1-\frac{1}{\alpha}\right)\in(0,1)$. Recursing from $t=0$ to $t=\tau^{\prime}-1$,
@@
Using Thm. 1 with $R=\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2}$ and setting $L_{0}=0$ completes the proof. ∎
@@
For an arbitrary comparator $u$ s.t. $f(u)\leq f(\theta_{T})$,
@@
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{1}{\lambda_{0}+\lambda_{1}f(\theta_{t})}[f(\theta_{t})-f(u)]$ (Using Lemma 1)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{1}{\lambda_{1}}\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}$ ( using $\lambda_{0}=0$ since $L_{0},\omega=0$ and by defining $C:=(2-\frac{1}{c})\,(1/\lambda_{1})$)
@@
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\alpha:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\alpha$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\alpha}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\alpha}\right)$
Since $f$ satisfies Assn. 1, 2 and 3 and $f(u)\leq\epsilon$, then, using Lemma 5 with $M=f(\theta_{0})$ and $B=\frac{e^{q}-1}{q}$ where $q$ is such that $\left\|\theta_{T}-u\right\|\leq\frac{q}{L_{1}}$. Since $\left\|\theta_{T}-u\right\|\leq\left\|\theta_{0}-u\right\|$, we can set $q=L_{1}\,\left\|\theta_{0}-u\right\|$. Hence,
$f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\underbrace{\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]}_{:=L}\,\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}$
$\implies f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\alpha}\right)\right]=\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\frac{\epsilon}{f(u)+\epsilon}\right]$ (19)
Hence, in order to ensure that $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\epsilon}{L}}{C}\,\left[1+\frac{f(u)}{\epsilon}\right].$ (Using the definition of $\alpha$)
@@
$T$ $\geq\frac{c\lambda_{1}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right].$
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Eq. 19 in Thm. 2 with the comparator $u=\beta u^{*}$ where $f(u)\leq\epsilon$. Hence, GD with Armijo line-search with $c$, $\eta_{\max}=\infty$ and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{2c-1}{c\lambda_{1}}$, $L:=\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]$ where $B=\frac{e^{q}-1}{q}$ and $q=L_{1}\,\left\|\theta_{0}-u\right\|$,
@@
$\implies f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{2}\right]$ (Since $f(\beta u^{*})\leq\epsilon$ and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$)
@@
$=\frac{216\,c\,}{(2c-1)(1-c)\gamma^{2}}\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$ (using Prop. 1 for the value of $\nu$ and $L_{1}$)